OpenStudy (anonymous):

what is the area of a circle if the radius is 6?

OpenStudy (anonymous):

Area of circle of radius "r" = pi . r^2 Area of circle of radius "6" = pi . 6^2 = 36 pi

OpenStudy (anonymous):

\[\huge \color{green}{Area = \pi r^2}\] Here, r = 6 Put: \[Area = \pi (6)^2 = \pi (36) = 36 \pi\] Put \(\pi = 3.14\) and then find the area..

OpenStudy (anonymous):

im watching you willie o.o

OpenStudy (anonymous):

A. 254.47 ft2 B. 27.28 ft2 C. 28.27 ft2 D. 63.62 ft2 here are the answers

OpenStudy (anonymous):

answer is c right ?

OpenStudy (anonymous):

Answer will be: 113.04 sq.feet..

OpenStudy (anonymous):

since we know that area of a circle is pi r^2 and here r =6 so, = (3.142) (6)^2 = 3.142 *36 =113.112

OpenStudy (anonymous):

113.112 m² = 1217.527 ft² but none of the answers matches this